# Statistical evaluation and optimization of entanglement purification protocols

### v1

@@ III.1 Markov chain Monte Carlo sampler
$\rho=\sum^{16}_{i=1}a_{i}B_{i}\quad\text{with}\quad\mathrm{Tr}\{\rho\}=1\quad\text{and}\quad\rho\geqslant 0,$ (38)
where $B_{i}$s are the basis vectors and $a_{i}\in\mathbb{R}$ for all $i$. The positive semidefinite condition $\rho\geqslant 0$ implies that $a_{i}$s have to fulfill three conditions based on Newton identities and Descartes’ rule of signs [32, 17]. Finally, one arrives at the result that all $4\times 4$ density matrices are presented by a $15$ dimensional convex body $K$ around the origin of $\mathbb{R}^{15}$, because due to $\mathrm{Tr}\{\rho\}=1$ one real parameter out of sixteen is fixed. Hence, if we consider $B_{16}=I_{4}/2$ then $a_{16}=1/2$ and the origin of $\mathbb{R}^{15}$ is the maximally mixed state. The vector ${\bf a}=(a_{1},a_{2},\dots a_{15})^{T}\in\mathbb{R}^{15}$ with the positive semidefinite condition yields a complete description of $K$. The hit-and-run algorithm introduced by Smith [33] realizes a random walk inside $K$ and it was shown that the underlying Markov chain converges to the uniform stationary distribution in polynomial time [34]. The convergence to the uniform distribution is independent of the starting point inside $K$ [35]. This algorithm provides a fast method of sampling large numbers ($10^{6}-10^{7}$) of density matrices. There exist other numerical approaches [17, 19, 36, 37, 20] that can also sample uniformly distributed density matrices, albeit with longer running times. All the statistical evaluations of the purification protocols are based on samples generated by the hit-and-run algorithm, whose details are shown in Appendix A.
Figure 1: (a) Average concurrence and (b) success probability as a function of the number of iterations for which a sample of $N=10^{7}$ density matrices was used (10 runs of the hit-and-run algorithm, each one outputting $10^{6}$ density matrices). Numerical evaluations are done for the four different purification protocols presented in Secs. II.1 – Bennett –, II.2 – Deutsch –, II.3 – MFI –, and II.4 – CNOT-based –. Horizontal dashed lines show the theoretical limits of the protocols for the average concurrence, which are given in Eqs. (53) – Bennett –, (54) – Deutsch –, (55) – MFI –, and (56) – CNOT-based –. The points are connected by lines to guide the eye. The MFI-based protocol and our newly proposed CNOT protocol turn out to produce the same average values, to the point that the green points are barely visible beyond the purple ones. The standard errors of the means are not visible in the plots. Values for the unbiased sample variance are available in Appendix B.
@@ III.2 Optimization with a quasi-Newton method
Our task is to improve the output state’s concurrence, which is a nonlinear function of $\boldsymbol{\alpha}_{AB}$, see Sec. II.5. In addition, based on Eq. (35) $\boldsymbol{\alpha}_{AB}$ has lower and upper bounds. Then, by using Eqs. (37) we define the cost function $f:\mathbb{R}^{30}\rightarrow\mathbb{R}$ as
@@
The cost function is based on the concurrence of the two-qubit state emerging after one iteration of the protocol. Let us consider that the sample of two-qubit density matrices consists $N$ elements $\rho_{1},...,\rho_{N}$, which are mapped to $\rho_{\text{out},1},...,\rho_{\text{out},N}$. The average output concurrence is estimated by
@@
which ensures that the optimal unitary transformations characterized by $\boldsymbol{\alpha}_{AB}$ increase the average concurrence of the whole sample. This method is the extension of our previous investigation [16], where the cost function depends only on $15$-dimensional vectors, and the optimal search is constrained to specific families of two-qubit states. We implement again one of the most effective quasi-Newton methods, the L-BFGS-B optimization algorithm [24, 25]. The gradient $\nabla\bar{f}(\boldsymbol{\alpha}_{AB})$ is obtained via automatic differentiation [27]. The algorithm does not require second derivatives, because the Hessian matrix is approximated. This approach yields a local minimum $\boldsymbol{\alpha}^{*}_{AB}$ of $\bar{f}$, i.e., $\bar{f}(\boldsymbol{\alpha}^{*}_{AB})\leqslant\bar{f}(\boldsymbol{\alpha}_{AB})$ for all $\boldsymbol{\alpha}_{AB}$ in the Euclidean norm defined neighborhood of $\boldsymbol{\alpha}^{*}_{AB}$.
Now, we briefly describe the methodology used to optimize the variational purification protocols of Sec. II.5. The optimization is subject to general two-qubit density matrices parametrized by a $15$-dimensional real vector $\boldsymbol{a}$. As the number of samples needed to cover the space of two-qubit quantum states with sufficient precision is particularly large ($10^{6}-10^{7}$ samples, see [38]), the algorithm struggles with a particularly slow optimization. Therefore, as a first attempt, we use smaller subsets of density matrices to find optimal unitary matrices and then test their performance on the whole sample.
@@
Input $\boldsymbol{\rho}(\boldsymbol{a}),\boldsymbol{a}\in\mathbb{R}^{15}$, ${P}_{1},\ {P}_{2},\ {P}_{3},\ {P}_{4}$ as in Eqs. (28), (29),  (30),  (31), $L_{\text{max}}$ iterations, optimizer (OPT)
@@ IV.1 Statistics of purification protocols
In this section, we compare the recurrence protocols of Secs. II.1, II.2, II.3, and II.4 based on their average performance on a sample of two-qubit states drawn from an almost uniform distribution [17]. More formally, let $\rho_{1},...,\rho_{N}$ be $N$ density matrices generated by the hit-and-run algorithm. As the composition of completely positive maps is again completely positive [39], every entanglement purification protocol presented in this work acts as a completely positive trace-preserving non-linear map. We denote this as $\Phi$, which maps the set of two-qubit quantum states
@@
where $P_{k}$ is the projector on $(A_{2},B_{2})$ defined in Eqs. (28), (29), (30), and (31). In the case of the protocols presented in Secs. II.1, II.2, II.3, different measurement results yield the same probability, and analytical formulas are given for the success probability $P^{(i)}_{s}(\rho_{j})$. The success probability of the approach in Sec. II.4 is also shown. However, in the optimized variational case, $P^{(i)}_{s}(\rho_{j})$ is a function of $P^{(i)}_{s}(k,\rho_{j})$, where $k$ is selected by the policy $\pi$ introduced in Eq. (LABEL:eq:pidefinition). It is worth noting that the condition in Eq.  (48) is decisive only in the first step for the protocols in Secs. II.1 and II.2, because after that all the remaining entangled two-qubit states are improved towards a Bell state. The average success probability is estimated after each iteration step by
@@
Figure 2: (a) Average concurrence and (b) success probability of the optimized purification protocols as a function of the number of iterations. The values shown in the plot represent the average over $N=10^{7}$ density matrices (10 runs of the hit-and-run algorithm, each one outputting $10^{6}$ density matrices). The asymptote of the MFI-based protocol in Eq. (55) and the ultimate limit of Eq. (57) are also shown. The points are connected by lines to guide the eye. The strategy of first destroying entanglement and then improving it (green points) delivers the best performance compared also to the protocols of Fig. 1. If we only keep the measurement result $k=1$ $(m=0,n=0)$, the average concurrence turns out to be lower compared to the case where the argmax policy of Eq. (34) is used. The standard error of the mean is not visible in the plot. Values for the unbiased sample variance are available in Appendix B.
@@
which represents the amount of entanglement present in the sample. Here we used parentheses to denote the standard error. We consider this number as a threshold and expect that entanglement purification protocols can improve it. The results of the four protocols are shown in Fig. 1. It is immediate to see that all protocols destroy entanglement in their very first iteration. The twirling operations used in both the Bennett and Deutsch protocol to obtain Werner or Bell diagonal states, respectively, destroy a significant portion of the entangled states. The remaining entangled states reach $\mathcal{C}=1$ in the limit of infinite iterations. We find these limits by imposing the conditions given by Eq. (7) and Eq. (11) on the whole sample. If a state does not satisfy it, then we assign $\mathcal{C}=0$ to it, otherwise, we set $\mathcal{C}=1$. In other words, every state satisfying the conditions can be purified into a Bell state in the limit of infinite iterations. This limit for the Bennett protocol yields
@@
while the Deutsch approach results
@@
which are based on the conditions in Eqs. (17), (18), (21), and (22). The performance of these protocols is almost the same and despite destroying entanglement in the first iteration, they can improve the average concurrence of the sample beyond its starting value. They achieve this already after four iterations (see Fig. 1). It is worth noting again that the asymptote of the Bennett protocol contains only the Bell state |$\ket{1}$, being a stable fixed point, while the other three protocols have two relevant stable fixed points in the limit of infinite iterations. These findings together with the average success probabilities characterize the performances and they show that both the MFI-based and the new CNOT protocols are superior to the pioneering ones. However, one can define an ultimate limit of every known or future proposal for entanglement purification, namely, when all entangled two-qubit states are converted to a maximally entangled state, which for our sample results in the numerical estimate
@@
Given this number, one can conclude that all four evaluated protocols are not very effective. The performance is confirmed by the analysis of the fidelities with respect to the stable fixed point states of the protocols – $r_{1}$ for Bennett, $r_{4}$ and $r_{2}$ for Deutsch, $r_{1}$ and $r_{2}$ for the MFI and $r_{4}$ and $r_{2}$ for the CNOT protocol – as a function of the number of protocol iterations, which are shown in Appendix C. This raises the question of how can one obtain better performance, which will be discussed in the subsequent section.
@@ IV.2 Optimization of variational recurrence protocols
In this section, we investigate numerically the variational purification protocol described in Sec. II. The optimization is based on the method given in Sec. III.2, where the gradient of the cost function associated with the protocol is computed through automatic differentiation. First, we sample $N=10^{6}$ density matrices using the hit-and-run algorithm. Afterwards, in order to speed up the optimization, we pick randomly $N_{s}=1000$ density matrices from this sample and calculate the gradient. This result in a numerical value for $\boldsymbol{\alpha}^{*}_{AB}$. We then apply the $\boldsymbol{\alpha}^{*}_{AB}$-dependent protocol on the whole sample. We also try out different strategies for the measurement policy $\pi$ of Eq. (LABEL:eq:pidefinition). In particular, we test the policy of Eq. (34) together with the case in which we only track the measurement result $k=1$ ($m=n=0$) in each iteration.
Our first try is the case when $\Pi$ in Eq. (27) is equal to $V(\boldsymbol{\alpha}_{AB})$ with policy $\pi$ of Eq. (34). In Fig. 2 we present the result, which demonstrates that this strategy does not destroy initial entanglement, as one expects. However, the actual increase in average concurrence is quite limited, which implies that an enormous amount of iterations are needed to purify the density matrices towards a maximally entangled state. As we seek to obtain a more efficient protocol, we need to consider additional operations performed on the sample of density matrices.
An interesting feature of Sec. IV.1 is that every protocol first destroys entanglement before it starts to improve the remaining two-qubit quantum states. Next, we consider the hypothesis that entanglement must be destroyed, to find an efficient protocol. Therefore, for the first iteration, we use the projector
@@
as the operation of the purification protocol, and for the rest of the iterations we make use of the optimization with $\Pi=V(\boldsymbol{\alpha}_{AB})$, which leads to higher average concurrence. Here, we have tested both of our measurement policies $\pi$. In both cases, these approaches outperform the limits of the MFI-based and the new CNOT protocols, which are shown in Fig. 2. We remind the reader that every optimized step of the variational protocol yields a different entangling gate. The resulting performance for many iterations seemingly approaches a limit, which is almost halfway between the ultimate asymptote and the asymptote defined by the MFI-based or the new CNOT protocol. The average success probability of this optimized protocol oscillates as a function of the number of iterations. We do not have a proper explanation for this effect, as it depends on the non-linear optimizer. However, we can see that the values of the success probability are almost always higher than those produced by the protocols considered in Fig. 1.
Finally, we need to address the realization of these optimized and abstract protocols. Concerning the two-qubit gates in $V(\boldsymbol{\alpha}_{AB})$, one can always employ quantum compilation [11, 40, 10], where the optimal unitary matrix $V(\boldsymbol{\alpha}^{*}_{AB})$ is translated into native gates on the chosen experimental platform. However, we are not aware of any possible implementation of $M_{2}$ in Eq. (58), but it seems necessary to first destroy entanglement before any variational purification protocol is applied.

### v2

@@ III.1 Markov chain Monte Carlo sampler
$\rho=\sum^{16}_{i=1}a_{i}B_{i}\quad\text{with}\quad\mathrm{Tr}\{\rho\}=1,\quad\rho\geqslant 0,$ (41)
where the $B_{i}$ are the basis vectors and $a_{i}\in\mathbb{R}$ for all $i$. The positive semidefinite condition $\rho\geqslant 0$ implies that the $a_{i}$ have to fulfill three conditions based on Newton identities and Descartes’ rule of signs [31, 17]. Finally, one arrives at the result that all $4\times 4$ density matrices are presented by a $15$ dimensional convex body $K$ around the origin of $\mathbb{R}^{15}$, because due to $\mathrm{Tr}\{\rho\}=1$ one real parameter out of sixteen is fixed. Hence, if we consider $B_{16}=I_{4}/2$ then $a_{16}=1/2$ and the origin of $\mathbb{R}^{15}$ is the maximally mixed state. The vector ${\bf a}=(a_{1},a_{2},\dots a_{15})^{T}\in\mathbb{R}^{15}$ with the positive semidefinite condition yields a complete description of $K$. The hit-and-run algorithm introduced by Smith [32] realizes a random walk inside $K$ and it was shown that the underlying Markov chain converges to the uniform stationary distribution in polynomial time [33]. The convergence to the uniform distribution is independent of the starting point inside $K$ [34]. This algorithm provides a fast method of sampling large numbers ($10^{6}-10^{7}$) of density matrices. There exist other numerical approaches [17, 19, 35, 36, 20] that can also sample uniformly distributed density matrices, albeit with longer running times. All the statistical evaluations of the purification protocols are based on samples generated by the hit-and-run algorithm, whose details are shown in Appendix A.
Figure 1: (a) Average concurrence and (b) success probability as a function of the number of iterations for which a sample of $N=10^{7}$ density matrices was used (ten runs of the hit-and-run algorithm, each one outputting $10^{6}$ density matrices). Numerical evaluations are done for the four different purification protocols presented in Secs. II.1II.4, i.e., Bennett, Deutsch, MFI-based, and CNOT-based protocols, respectively. Horizontal dashed lines show the theoretical limits of the protocols for the average concurrence, which are given in Eqs. (56)–(59) for the Bennett, Deutsch, MFI-based, and CNOT-based protocols, respectively. The points are connected by lines to guide the eye. The MFI-based protocol and our proposed CNOT protocol turn out to produce the same average values, to the point that the green points are barely visible beyond the purple ones. The standard errors of the means are not visible in the plots. Values for the unbiased sample variance are available in Appendix B.
@@ III.2 Optimization with a quasi-Newton method
Our task is to improve the concurrence of the output state, which is a nonlinear function of $\boldsymbol{\alpha}_{AB}$ (see Sec. II.5). In addition, based on Eq. (30) $\boldsymbol{\alpha}_{AB}$ has lower and upper bounds. Then, by using Eq. (40) we define the cost function $f:\mathbb{R}^{30}\rightarrow\mathbb{R}$ as
@@
The cost function is based on the concurrence of the two-qubit state emerging after one iteration of the protocol. Let us consider that the sample of two-qubit density matrices consists of $N$ elements $\rho_{1},...,\rho_{N}$, which are mapped onto $\rho_{\text{out},1},...,\rho_{\text{out},N}$. The average output concurrence is estimated by
@@
which ensures that the optimal unitary transformations characterized by $\boldsymbol{\alpha}_{AB}$ increase the average concurrence of the whole sample. This method is the extension of the one developed in our previous investigation [16], where the cost function depends only on $15$-dimensional vectors, and the optimal search is constrained to specific families of two-qubit states. We implement again one of the most effective quasi-Newton methods, the L-BFGS-B optimization algorithm [23, 24]. The gradient $\nabla\bar{f}(\boldsymbol{\alpha}_{AB})$ is obtained via automatic differentiation [26]. The algorithm does not require second derivatives, because the Hessian matrix is approximated. This approach yields a local minimum $\boldsymbol{\alpha}^{*}_{AB}$ of $\bar{f}$, i.e., $\bar{f}(\boldsymbol{\alpha}^{*}_{AB})\leqslant\bar{f}(\boldsymbol{\alpha}_{AB})$ for all $\boldsymbol{\alpha}_{AB}$ in the Euclidean norm defined neighborhood of $\boldsymbol{\alpha}^{*}_{AB}$. Now, we briefly describe the methodology used to optimize the variational purification protocols of Sec. II.5. The optimization is subject to general two-qubit density matrices parametrized by a $15$-dimensional real vector $\boldsymbol{a}$. As the number of samples needed to cover the space of two-qubit quantum states with sufficient precision is particularly large ($10^{6}-10^{7}$ samples; see [37]), the algorithm struggles with a particularly slow optimization. Therefore, as a first attempt, we use smaller subsets of density matrices to find optimal unitary matrices and then test their performance on the whole sample.
@@
Input $\boldsymbol{\rho}(\boldsymbol{a}),\boldsymbol{a}\in\mathbb{R}^{15}$, ${P}_{1},\ {P}_{2},\ {P}_{3},\ {P}_{4}$ as in Eqs. (33), (34),  (35), and (36), $L_{\text{max}}$ iterations, optimizer (OPT)
@@ IV.1 Statistics of purification protocols
In this section, we compare the recurrence protocols of Secs. II.1, II.2, II.3, and II.4 based on their average performance on a sample of two-qubit states drawn from an almost uniform distribution [17]. More formally, let $\rho_{1},...,\rho_{N}$ be $N$ density matrices generated by the hit-and-run algorithm. As the composition of completely positive maps is again completely positive [38], every entanglement purification protocol presented in this work acts as a completely positive trace-preserving non-linear map. We denote this by $\Phi$, which maps the set of two-qubit quantum states
@@
where $P_{k}$ is the projector on $(A_{2},B_{2})$ defined in Eqs. (33), (34), (35), and (36). In the case of the protocols presented in Secs. II.1, II.2, II.3, different measurement results yield the same probability, and analytical formulas are given for the success probability $P^{(i)}_{s}(\rho_{j})$. In the case of the MFI-based protocol, the success probability is given by the probability of successfully performing the quantum operation given in Eq. (13) and the probability of measuring the state in one of the four possible outcomes is $1/4$. The success probability of the approach in Sec. II.4 is also shown. However, in the optimized variational case, $P^{(i)}_{s}(\rho_{j})$ is a function of $P^{(i)}_{s}(k,\rho_{j})$, where $k$ is selected by the policy $\pi$ introduced in Eq. (LABEL:eq:pidefinition). It is worth noting that the condition in Eq.  (51) is decisive only in the first step for the protocols in Secs. II.1 and II.2, because after that all the remaining entangled two-qubit states are improved towards a Bell state. The average success probability is estimated after each iteration step by
@@
Figure 2: (a) Average concurrence and (b) success probability of the optimized purification protocols as a function of the number of iterations. The values shown in the plot represent the average over $N=10^{7}$ density matrices (ten runs of the hit-and-run algorithm, each one outputting $10^{6}$ density matrices). The asymptote of the MFI-based protocol in Eq. (58) and the ultimate limit of Eq. (60) are also shown. The points are connected by lines to guide the eye. The strategy of first destroying entanglement and then improving it (green points) delivers the best performance compared also to the protocols of Fig. 1. If we only keep the measurement result $k=1$ $(m=0,n=0)$, the average concurrence turns out to be lower compared to the case where the argmax policy of Eq. (39) is used. The standard error of the mean is not visible in the plot. Values for the unbiased sample variance are available in Appendix B.
@@
which represents the amount of entanglement present in the sample. Here we use parentheses to denote the standard error. We consider this number as a threshold and expect that entanglement purification protocols can improve it. The results of the four protocols are shown in Fig. 1. It is immediate to see that all protocols destroy entanglement in their very first iteration. The twirling operations used in both the Bennett and Deutsch protocols to obtain Werner or Bell diagonal states, respectively, destroy a significant portion of the entangled states. The remaining entangled states reach $\mathcal{C}=1$ in the limit of infinite iterations. We find these limits by imposing the conditions given by Eqs. (8) and (12) on the whole sample. If a state does not satisfy it, then we assign $\mathcal{C}=0$ to it, otherwise, we set $\mathcal{C}=1$. In other words, every state satisfying the conditions can be purified into a Bell state in the limit of infinite iterations. This limit for the Bennett protocol yields
@@
while the Deutsch approach results in
@@
which are based on the conditions in Eqs. (18), (19), (23), and (24). The performance of these protocols is almost the same and despite destroying entanglement in the first iteration, they can improve the average concurrence of the sample beyond its starting value. They achieve this already after four iterations (see Fig. 1). It is worth noting again that the asymptote of the Bennett protocol contains only the Bell state |$\ket{1}$, being a stable fixed point, while the other three protocols have two relevant stable fixed points in the limit of infinite iterations. These findings together with the average success probabilities characterize the performances and they show that both the MFI-based and our CNOT protocols are superior to the pioneering ones. However, one can define an ultimate limit of every known or future proposal for entanglement purification, namely, when all entangled two-qubit states are converted to a maximally entangled state, which for our sample results in the numerical estimate
@@
Given this number, one can conclude that all four evaluated protocols are not very effective. The performance is confirmed by the analysis of the fidelities with respect to the stable fixed-point states of the protocols, i.e., $r_{1}$ for the Bennett, $r_{4}$ and $r_{2}$ for the Deutsch, $r_{1}$ and $r_{2}$ for the MFI and $r_{4}$ and $r_{2}$ for the CNOT protocols, as a function of the number of protocol iterations, which are shown in Appendix C. This raises the question of how one can obtain better performance, which will be discussed in the subsequent section.
@@ IV.2 Optimization of variational recurrence protocols
In this section, we investigate numerically the variational purification protocol described in Sec. II. The optimization is based on the method given in Sec. III.2, where the gradient of the cost function associated with the protocol is computed through automatic differentiation. First, we generate $N=10^{6}$ density matrices using the hit-and-run algorithm to obtain an almost uniformly distributed sample. Afterward, to speed up the optimization, we randomly pick $N_{s}=1000$ density matrices from this sample and calculate the gradient. This approach guarantees that optimization is less time consuming and the convex set of all two-qubit states is well represented. Nevertheless, the sampling of only $10^{3}$ density matrices is not enough to ensure uniform distribution, in particular for the statistics of the CNOT and MFI protocols. The optimization results in a numerical value for $\boldsymbol{\alpha}^{*}_{AB}$. We then apply the $\boldsymbol{\alpha}^{*}_{AB}$-dependent protocol on the whole sample. We also try different strategies for the measurement policy $\pi$ of Eq. (LABEL:eq:pidefinition). In particular, we test the policy of Eq. (39) together with the case in which we only track the measurement result $k=1$ ($m=n=0$) in each iteration.
Our first try is the case when $\Pi$ in Eq. (32) is equal to $V(\boldsymbol{\alpha}_{AB})$ with policy $\pi$ of Eq. (39). In Fig. 2 we present the result, which demonstrates that this strategy does not destroy initial entanglement, as one expects. However, the actual increase in average concurrence is quite limited, which implies that an enormous number of iterations are needed to purify the density matrices towards a maximally entangled state. As we seek to obtain a more efficient protocol, we need to consider additional operations performed on the sample of density matrices.
An interesting feature of Sec. IV.1 is that every protocol first destroys entanglement before it starts to improve the remaining two-qubit quantum states. Next, we consider the hypothesis that entanglement must be destroyed in order to find an efficient protocol. Therefore, for the first iteration, we use the projector
@@
as the operation of the purification protocol, and for the rest of the iterations we make use of the optimization with $\Pi=V(\boldsymbol{\alpha}_{AB})$, which leads to higher average concurrence. Here, we have tested both of our measurement policies $\pi$. In both cases, these approaches outperform the limits of the MFI-based and the CNOT protocols, which are shown in Fig. 2. We remind the reader that every optimized step of the variational protocol yields a different entangling gate. The resulting performance for many iterations seemingly approaches a limit, which is almost halfway between the ultimate asymptote and the asymptote defined by the MFI-based or the CNOT protocol. The average success probability of this optimized protocol oscillates as a function of the number of iterations. We do not have a proper explanation for this effect, as it depends on the non-linear optimizer. However, we can see that the values of the success probability are almost always higher than those produced by the protocols considered in Fig. 1.
Finally, we need to address the realization of these optimized and abstract protocols. Concerning the two-qubit gates in $V(\boldsymbol{\alpha}_{AB})$, one can always employ quantum compilation strategies [11, 39, 40, 10], where the optimal unitary matrix $V(\boldsymbol{\alpha}^{*}_{AB})$ is translated into native gates on the chosen experimental platform. However, we are not aware of any possible implementation of $M_{2}$ in Eq. (61), but it seems necessary to first destroy entanglement before any variational purification protocol is applied.